OpenStudy (scarlettfarra2000):

Please help me, Solve the equation below. Don't forget to first simplify each side of the equation, if possible check each solution

OpenStudy (scarlettfarra2000):

\[\frac{ 1 }{ 5 }x-3=-\frac{ 4 }{ 5 }x-3\]

OpenStudy (scarlettfarra2000):

We need to get x alone on one side of the equation and all of the numbers on the other side

OpenStudy (scarlettfarra2000):

I was thinking of getting x-term together on the left side of the equation, and then simply. I added \[\frac{ 1 }{ 5 }x\] to both sides and combine like terms

OpenStudy (anonymous):

i hate fractions !

OpenStudy (scarlettfarra2000):

\[\frac{ 1 }{ 5 }x-3+\frac{ 1 }{ 5 }x=-\frac{ 4 }{ 5 }x-3+\frac{ 1 }{ 5 }x\]

OpenStudy (anonymous):

multiply both sides by 5, and you will get rid of them

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 }x-3=-\frac{ 4 }{ 5 }x-3\\ 5\left(\frac{ 1 }{ 5 }x-3\right)=5\left(-\frac{ 4 }{ 5 }x-3\right)\]

OpenStudy (anonymous):

or if you want to do it the easy way, add \(3\) to both sides first! that will give you \[\frac{1}{5}x=-\frac{4}{5}x\]